what is vant hoff factor for nacl
what is vant hoff factor for alcl3
 

Grade:12th pass

3 Answers

Akshat Agarwal
7 years ago
Vont hoff favtor of alcl3 is 1 because 
Because alcl3 is covalent molecule. So n=1
So by i=1+a(n-1)
Vont hoff factor is1
ADARSH SANKAR R S
7 years ago
Vant Hoff factor of AlCl3 is 1.
Reason:AlCl3 is a covalent molecule....
Formula: i=1+a(n-1)..
Thank you..
 
Umakant biswal
7 years ago
DEAR JHEEL
For ions with a one to one ratio, like NaCl, this dissociates into ions of 
Na+1 and Cl-1 and the van't Hoff factor is then 2. 
in the same way , alcl3 will dissociate into al^+3 and 3cl- , so, total n = 4 
i = 1+1 (4-1) 
i = 4 
HOPE IT CLEARS NOW 
REGARDS